OpenStudy (anonymous):

Need Alebra help please, the person helping me before had to go. I've been stuck on this assigment for 3 hours now. I already have parts of it done. So I need to Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real. So far I go 0^2 0 4(1) c = -1. All I need is to figure out c and try to solve the problem

OpenStudy (anonymous):

All I need is to know what number I should put for c that would equal -1 and solve it, please help

OpenStudy (anonymous):

I mean 0^2+4(1)c = -1

OpenStudy (cwrw238):

oh ok so you need to solve 0 - 4c = -1 c = 1/4 so b = 0, a = 1 and c = 1/4 giving the equation 1x^2 + 0x - 1/4 = 0 x^2 + 1/4 =

OpenStudy (cwrw238):

x^2 + 1/4 = 0

OpenStudy (anonymous):

is that all the steps you have to do?

OpenStudy (cwrw238):

i'm not entirely sure what you want here you have b = 0 but it could be = 2 say the discriminant is b^2 - 4ac where a, b and c are the constants in the quadratic equation a^2 + bx + c = 0 the sign of the discriminant tells you whether the roots are real and different, real and equal or imaginary. if discriminant is negative they are imaginary because there is no real square root of a negative

OpenStudy (anonymous):

I think that this is what she's looknig for! EIther way I'm unable to continue to be on openstudy I gtg. Thanks so much for your help!

OpenStudy (cwrw238):

to find the solution of x^2 + 1/4 = 0 we can use the formula for solution of quadratic x = [-b +/- sqrt( b^2 - 4ac) ]/ 2a here a = 1, b = 0 and c = 1/4 so x = [-0 +/- sqrt(0 - 4*1*(1/4)] / 2 = +/- sqrt(-1) / 2 = + (½)i , - (½)i where i = sort (-1)

OpenStudy (cwrw238):

* sqrt -1

OpenStudy (anonymous):

Thanks again :) I really appreciate it. I wish I could give you more medals.

OpenStudy (cwrw238):

hope this helps